# The Sound of Silence in Social Networks

### v1

@@ 3.2 Consensus in $\mbox{SOM}^{-}$  Cliques
Assume, by contradiction, that from some time $t$ onwards, no agent with an opinion value greater or equal to $U$ will speak; this implies that all the agents that speak at time $t$ and after time $t$ will have opinion values lower than $U$. In fact, from Eq. 2, we know that the highest opinion of any non-silent agent from time $t$ onwards will be at most $max({\mathbf{B}^{t}_{i}\ |\ \mathbf{B}^{t}_{i}<U\ \mbox{for any}\ i\in A})$.
From Cor. 1 and the contradiction assumption, at least one of these non-silent agents, whose opinion is less than or equal to $max({\mathbf{B}^{t}_{i}\ |\ \mathbf{B}^{t}_{i}<U\ \mbox{for any}\ i\in A})$, will speak infinitely often.
Since the graph is a clique, the opinion values of all agents with opinion values greater than or equal to $U$ at time $t$ will decrease, moving, progressively, closer to a value lower or equal to $max({\mathbf{B}^{t}_{i}\ |\ \mathbf{B}^{t}_{i}<U\ \mbox{for any}\ i\in A})$; therefore, the opinion values of these agents will not converge to $U$; which contradicts Th. 3.1.
From the above, to prove consensus (i.e., $U=L$), we will now show how bounded the distance between the maximum and minimum opinion values is at different time units.
For the following lemma, we need to consider the minimum influence of the graph, defined as $I_{min}=\min_{(i,j)\in E}I(i,j)$ and the difference between the maximum and minimum opinion value at time $t$, defined as $R_{t}=max(\mathbf{B}^{t})-\min(\mathbf{B}^{t})$, notice that $R_{t}\in[0,1]$ at any time $t$.
Lemma 4 ($m-\epsilon$ decrement)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with an $n$-agent influence graph $G=(A,E,I)$ where $G$ is a clique with $n\geq 3$. For all $m\in\mathbb{N}$, there exists a $t\in\mathbb{N}$ such that
$R_{0}-R_{t}\geq m*\epsilon$
with $\epsilon=I_{min}\cdot(U-L)$.
@@
It can be proved by induction on m for all $m\in\mathbb{N}$ by using Lem. 3; notice that in each time unit $t$ where there is at least a non-silent agent with an opinion greater or equal to $U$, the distance between the maximum and minimum opinions will be reduced by a value greater or equal to a constant factor, specifically $R_{t+1}-R_{t}\geq\epsilon$ with $\epsilon=I_{min}\cdot(U-L)$.
Now, we can state our consensus result for cliques with at least three agents.
Theorem 3.2 (Consensus in $\mbox{SOM}^{-}$ Cliques)
Proof
From Th. 3.1, there exists $U,\,L\in[0,1]$ such that $U=\lim_{t\to\infty}max(\mathbf{B}^{t})$ and $L=\lim_{t\to\infty}min(\mathbf{B}^{t})$.
We now prove $U=L$ by contradiction.
@@
Suppose, by contradiction, that $U\neq L$. As $U$ can not be lower than $L$, therefore, $U-L>0$. As $I_{min}>0$, $\epsilon$ (i.e. $I_{min}\cdot(U-L)$) is a constant greater than zero. Thus, there exist some $m\in\mathbb{N}$ such that $m*\epsilon>1$, however, from Lem. 4 there must exist a time $t$ where $R_{0}-R_{t}\geq m*\epsilon>1$, therefore, $R_{0}>1+R_{t}$, but it is not possible as $R_{0},\ R_{t}\in[0,1]$, which is a contradiction.
Therefore, $U=L$ and the model converges to consensus.
Remark 3
Notice that for $2$-agent cliques, consensus is not guaranteed; let us consider a clique with two agents: agent $1$ and agent $2$, where the opinions are $\mathbf{B}^{0}_{1}=1$ and $\mathbf{B}^{0}_{2}=0$, the influences are $I(1,2)=I(2,1)=1$ and the tolerance radii are $\tau_{1}=\tau_{2}=1$. Suppose that $\mathbf{B}^{0}=(1,1).$
In this case, the opinion evolution of agent $1$, starting with opinion $1$, and agent $2$, starting with opinion 00, always alternate between the values $1$ and 00. Figure 1 illustrates the opinion evolution of this clique.
(a) Influence graph
(b) State evolution
Figure 1: Opinion evolution in a $2-$agent clique with influence 1, with $\mathbf{B}^{0}=(1,0)$.
3.3 Dissensus in $\mbox{SOM}^{-}$ Models
In strongly connected aperiodic graphs, the $\mbox{SOM}^{-}$ model differs from the DeGroot model by no longer guaranteeing consensus. Agents can enter a state of perpetual silence, effectively disrupting opinion propagation as if severing connections from the graph. This phenomenon is particularly critical when silent agents form bridges between connected components. Their opinions, influenced by opposing connected components, may remain below the 50% non-silent threshold indefinitely. As a result, they prevent opinion exchange between components, obstructing the possibility of achieving consensus. The following example illustrates this scenario through by showing the agents state evolution and influence graph (see Figure 2).
Remark 4
For visual clarity, self-influences are omitted from the influence graphs. As a result, the visible incoming influences for each agent may not sum to one. Nevertheless, self-influences are implicitly present to ensure conformity with Definition 1. Readers can infer an agent’s self-influence by subtracting the sum of its visible incoming influences from one.
(b) Each plot shows agents’ state evolution over time. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.8,0.5,0.2,0.0)$.
Figure 2: Examples of Dissensus in $\mbox{SOM}^{-}$ Models
In this example(Figure 2), agent $3$ acts as a critical bridge between two graph components: agents $1$ and $2$, and agents $4$ and $5$. Initially, all agents have enough opinions within their tolerance radii, except for agent $3$’s. This causes only agent $3$ to become silent at $t=1$, effectively severing the only link between the two components. As the components converge to their local consensus values, agent $3$’s opinion remains static due to equal influences from both components. These evolving component opinions consistently fall outside agent $3$’s tolerance radius, trapping it in perpetual silence. Consequently, the two components remain permanently disconnected, preventing global consensus and exemplifying how the $\mbox{SOM}^{-}$ model can lead to dissensus in strongly connected graphs.
3.4 Dissensus in $\mbox{SOM}^{+}$ Models
For the $\mbox{SOM}^{+}$ models, consensus cannot even be guaranteed for clique graphs. While this model shares similarities with $\mbox{SOM}^{-}$ regarding perpetual silence, it introduces a unique phenomenon where the entire graph can enter and indefinitely remain in a silent state. This distinction stems from how silent agents influence the opinion update in $\mbox{SOM}^{+}$.
Recall from Section 2.3.2 that $\text{p}\mathbf{B}^{t}$ is the public state of opinion at time $t$ and represents the most recent public opinion of each agent. Henceforth we will refer to $\text{p}\mathbf{B}^{t}_{i}$ as the public opinion of agent $i$.
In $\mbox{SOM}^{+}$, non-silent agents influence the opinions of their neighbors as in $\mbox{SOM}^{-}$. But, silent agents influence the opinions of their neighbors with their public opinion instead of excluding their opinion from the updates as in $\mbox{SOM}^{-}$.
Consequently, if an agent’s opinion converges to a value that places the last public opinions of more than half of its neighbors outside its tolerance range, the agent will remain perpetually silent unless its neighbors change their public opinions. This scenario can lead to a state where intuitively some or all agents become “tire” of the discourse and cease participation (i.e., become silent), while the remaining active agents continue to consider the public opinions of these silent agents.
Unlike $\mbox{SOM}^{-}$, where opinions of silent agents are disregarded, $\mbox{SOM}^{+}$ allows for the persistence of unchanged public opinions indefinitely. This characteristic can result in dissensus due to the formation of public opinions that no longer reflect the current opinions of silent agents.
To illustrate these dynamics, we present a dissensus example (Figure 3) that with a clique influence graph where all agents initially have more than half of their neighbors’ opinions outside their tolerance radius. Consequently, all agents become silent at $t=1$. At this point, each agent’s updated opinion places all other agents’ opinions outside its tolerance radius. The agents’ opinions then converge to distinct values, determined by the initial opinions $\mathbf{B}^{0}$, which perpetuates the condition for silence. This state persists indefinitely, as the convergence values maintain the silence condition for all agents. The result is dissensus. This example demonstrates how the $\mbox{SOM}^{+}$ model can lead to complete silence and opinion divergence in a clique, preventing the possibility of consensus.
(a) Clique influence graph (self influence excluded)
(b) Each plot shows agents’ state evolution over time. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$.
Figure 3: Example of Dissensus in $\mbox{SOM}^{+}$ Model
4 Experimenting with the spiral of silence
To explore the implications of our proposed models, we developed a simulator, which generated all the examples presented in this paper. This section illustrates specific scenarios providing insights into opinion dynamics under the spiral of silence.
We begin by demonstrating how our $\mbox{SOM}^{-}$ model reflects the phenomenon of reinforcement of dominant views, a key aspect of the spiral of silence theory introduced earlier. This example illustrates, as predicted by the theory, how a vocal minority can disproportionately influence overall opinion in a network despite being outnumbered.
Next, we present a noteworthy case from our $\mbox{SOM}^{+}$ model where agents’ opinions converge to consensus while their public opinions remain divergent. This scenario mirrors situations in social media where individuals may unknowingly share common views, yet perceive a state of disagreement due to the persistence of outdated public opinions as discussed [17].
Finally, we outline our current experimental capabilities, showcasing how our simulator can model opinion evolution in networks that more closely approximate real-world social structures in both topology and scale.
4.1 Reinforcement of Dominant Views
Figure 4 illustrates how our $\mbox{SOM}^{-}$ model captures the essence of the spiral of silence theory. In this strongly connected influence graph, a vocal minority effectively dominates the discourse, causing the silent majority to converge towards the perceived majority opinion.
The network comprises two groups: Group 1 (Agents 1, 2, 3, 6, 7, and 8) with opinions in the lower half of the spectrum, and Group 2 (Agents 4 and 5) with opposing views near the higher end. Despite Group 1 being the actual majority, the network topology and differing tolerance radii $\tau$ lead to Group 2 dominating the opinion dynamics. Group 1 agents lack intra-group connections but are all linked to Group 2 agents. Moreover, Group 2’s significantly larger tolerance radii allow them to remain non-silent more frequently.
This configuration results in the vocal minority (Group 2) disproportionately influencing the network. The silent majority (Group 1) remains quiet for most of the update process, only becoming active when opinions have already shifted closer to the perceived majority view. This example demonstrates how the $\mbox{SOM}^{-}$ model can simulate scenarios where a minority opinion, through strategic positioning and persistent vocalization, can shape the overall opinion landscape, even when numerically outnumbered.
(b) Each plot shows agents’ state evolution over time. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(0.1,0.2,0.15,1.0,0.85,0.25,0.3,0.05)$ and $\tau=(0.1,0.05,0.1,0.85,0.6,0.05,0.1,0.05)$.
Figure 4: Silent Majority vs. Vocal Minority: Opinion Dynamics in $\mbox{SOM}^{-}$
4.2 Hidden Consensus in $\mbox{SOM}^{+}$ Model
The $\mbox{SOM}^{+}$ model reveals a noteworthy phenomenon: the possibility of reaching a consensus that remains undetected by the agents themselves. Figure 5 illustrates this scenario using a clique graph with four agents.
Initially, the agents hold diverse opinions ($\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$). However, due to the graph’s influence structure, all agents become silent after $t=0$. Despite this silence, their private opinions converge to a common value (approximately 0.5) over time. Crucially, this convergence occurs without any further public expression of opinions, leaving each agent unaware of the emerging consensus.
This hidden consensus phenomenon mirrors real-world scenarios in social media where individuals may unknowingly share common views while perceiving disagreement due to outdated public expressions [17].
(a) Influence graph self influence excluded
(b) Each plot shows agents’ state evolution over time. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$.
Figure 5: Silent Consensus Scenario
4.3 Scaling to Large-Scale Network Simulations
While the small-scale examples presented earlier are valuable for demonstrating mathematical properties, they fail to capture the complexity of real-world social networks. To address this limitation, we developed a simulation platform capable of modeling networks with over a million agents, providing insights into how our models behave in more realistic scenarios.
Our approach utilizes a highly optimized version of the preferential attachment algorithm [1] to generate networks exhibiting small-world properties [25] and power-law degree distributions. These networks are commonly for experimentation because they reflect real-world social networks. To manage the computational demands of large-scale simulations, we employ parallel processing via Scala and Akka Actors [19, 18], with results stored in a PostgreSQL database for efficient analysis.
Simulations of these large-scale networks reveal intriguing dynamics in both the $\mbox{SOM}^{+}$ and $\mbox{SOM}^{-}$ models. In the $\mbox{SOM}^{+}$ model, we observe that increased network connectivity hinders consensus formation. Conversely, less dense networks facilitate consensus more readily. The $\mbox{SOM}^{-}$ model exhibits opposite behavior: higher connectivity promotes the formation of a global spiral of silence, driving silent agents’ beliefs towards the perceived majority opinion. In contrast, sparser networks tend to foster local spirals of silence or echo chambers, often interconnected by "bridge" agents that may become perpetually silent, impeding global consensus.
These findings align with previous research [23, 7], which similarly identified the emergence of local spirals of silence in lower-density networks and global spirals of silence in denser networks. Our large-scale simulations provide a computational framework for exploring the intricate relationship between network structure and opinion dynamics under the influence of the spiral of silence.
5 Conclusions and Related Work
In this work, we have extended the classical DeGroot model to incorporate key social dynamics described by the Spiral of Silence. Our contributions highlight how the addition of memory, even in a limited form as in $\mbox{SOM}^{+}$, introduces significant complexity to consensus-building processes. Through theoretical results and simulations, we demonstrated that consensus, while achievable in fully connected networks (i.e., cliques) in $\mbox{SOM}^{-}$, is no longer guaranteed, in contrast with the DeGroot model, in more general strongly-connected aperiodic graphs. This points to the nuanced impact that silence and memory can have on opinion formation in social networks. It also offers insights into the challenges of converging to consensus in real-world scenarios. We also discuss simulations reflecting predictions of the Spiral of Silence such as the reinforcement of dominant views and hidden consensus. Finally we discussed the simulation in our models of large-scale networks.
We are not aware of any prior work that extends DeGroot-based models to incorporate the Spiral of Silence. However, some recent studies have explored the Spiral of Silence theory within agent-based networks. In [23], the authors analyzed the impact of manipulative actors in social networks by building an agent-based model grounded in Spiral of Silence theory and complex adaptive systems. In this model, agents hold a binary opinion (agree/disagree) that remains fixed over time but may choose whether or not to express it, depending on the prevailing opinion climate. In [7], authors investigate how the number of communities in a network and connectivity between them affects the perceived opinion climate. Nevertheless, these works do not deal with convergence to consensus, opinion updates or memory as done in this paper.
The exclusion of silent agents in $\mbox{SOM}^{-}$ amounts to having edges (influences) disappearing and reappearing during opinion evolution which reflects the dynamic influence nature of this model. There are several works studying dynamic influence in opinion formation. The work [13] introduces a version of the DeGroot model in which self-influence changes over time while the influence on others remains the same. The works [9, 10] explore convergence and stability, respectively, in models where influences change over time. The work [6] demonstrates how asynchronous communication, when combined with dynamic influence, can prevent consensus. None of this work deals with the dynamics derived from the Spiral of Silence.
Funding
This work is supported by the Colombian Minciencias project PROMUEVA, BPIN 2021000100160.

### v2

@@ Appendix 0.A Appendix
First, we prove that at each time unit $t$ where there is at least one non-silent agent, the distance between the maximum and minimum opinions will be reduced by at least a factor of $(1-I_{min})$ at time $t+1$, i.e. $R_{t+1}\leq(1-I_{min})\cdot R_{t}$ where $1-I_{min}$ is a positive constant below one.
We consider the following two cases for each unit time $t$ where there is at least one non-silent agent:
  1. 1.
    At least one agent is silent. In this case, we consider a non-silent agent $k$, the updated opinion of every agent $i$ influenced by agent $k$ will have an upper bound, thus:
    $\mathbf{B}^{t+1}_{i}=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    And the lower bound would be the following:
    $\mathbf{B}^{t+1}_{i}=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    Therefore:
    $(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}\geq\mathbf{B}^{t+1}_{i}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$ (7)
    As the graph is a clique and there is at least one silent agent, every agent in the graph is influenced by agent $k$, including itself; notice that as there is at least one silent agent, call it $l$, then $(1-\sum_{j\in N_{i}}I_{ji})$ from Eq. 4, corresponding to the self-influence on agent $i$, is at least $I_{li}\geq I_{min}$. Therefore, all agents in the graph satisfy Eq. 7; then, there are both an upper and lower bound for all opinions in unit time $t+1$, including the maximum and the minimum opinions, thus:
    $\max(\mathbf{B}^{t+1})\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$ (8)
    $\min(\mathbf{B}^{t+1})\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$ (9)
    Then, as $R_{t+1}=\max(\mathbf{B}^{t+1})-\min(\mathbf{B}^{t+1})$, $R_{t+1}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}-(1-I_{min})\cdot\min(\mathbf{B}^{t})-I_{min}\cdot\mathbf{B}^{t}_{k}=(1-I_{min})\cdot(\max(\mathbf{B}^{t})-\min(\mathbf{B}^{t}))=(1-I_{min})\cdot R_{t}$. Then $R_{t+1}\leq(1-I_{min})\cdot R_{t}$.
  2. 2.
    All agents speak at time $t$: as the graph is a clique with $n\geq 3$, there must be at least one non-silent agent $k$ that is not the unique maximum or minimum opinion agent in the graph  [^4] Notice that this case is not applicable in the clique in Remark 3, as such a clique does not have a non-silent agent that is not the unique maximum or minimum opinion agent in the graph.; agent $k$ influences all the other agents such that their opinions are bounded as follows:
    $\mathbf{B}^{t+1}_{i}=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    $\mathbf{B}^{t+1}_{i}=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    As agent $k$ is influenced by an agent with maximum opinion and an agent with minimum opinion at time $t$, its opinion at time $t+1$ will be bounded as follows:
    $\mathbf{B}^{t+1}_{k}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\max(\mathbf{B}^{t})$
    $\mathbf{B}^{t+1}_{k}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\min(\mathbf{B}^{t})$
    Therefore:
    $\mathbf{B}^{t+1}_{k}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\max(\mathbf{B}^{t})\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    and
    $\mathbf{B}^{t+1}_{k}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\min(\mathbf{B}^{t})\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$
    Therefore, every agent in the graph, including $k$, holds the following bounds:
    $(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}\geq\mathbf{B}^{t+1}_{i}\geq(1-I_{min})\cdot\min(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}$ (10)
    Then, as $R_{t+1}=\max(\mathbf{B}^{t+1})-\min(\mathbf{B}^{t+1})$, $R_{t+1}\leq(1-I_{min})\cdot\max(\mathbf{B}^{t})+I_{min}\cdot\mathbf{B}^{t}_{k}-(1-I_{min})\cdot\min(\mathbf{B}^{t})-I_{min}\cdot\mathbf{B}^{t}_{k}=(1-I_{min})\cdot(\max(\mathbf{B}^{t})-\min(\mathbf{B}^{t}))=(1-I_{min})\cdot R_{t}$. Then $R_{t+1}\leq(1-I_{min})\cdot R_{t}$.
Thus, we have proved that $R_{t+1}\leq(1-I_{min})\cdot R_{t}$ for any time unit $t$ where there is at least one non-silent agent.
Finally, by using the above and Cor. 1, it can be proved, by induction, that for all $m\in\mathbb{N}$, there exists $t\in\mathbb{N}$ such that $R_{t}\leq R_{0}\cdot(1-I_{min})^{m}$.
See 3.2
@@
From Th. 3.1, there exist $U,\,L\in[0,1]$ such that $U=\lim_{t\to\infty}max(\mathbf{B}^{t})$ and $L=\lim_{t\to\infty}min(\mathbf{B}^{t})$. We now prove $U=L$ by contradiction. Suppose that $U\neq L$. As $U$ can not be lower than $L$, therefore, UL>0U-L>0. From $\lim_{m\to\infty}R_{0}\cdot(1-I_{min})^{m}=0$  [^5] Notice that $R_{0}$ and $I_{min}$ are positive constants such that $1>(1-I_{min})>0$. and Lem. 3 there must exists a time $t$ where $R_{t}<(U-L)$, but it is not possible as, from Th. 3.1 and Cor.  2, $\min(\mathbf{B}^{t})\leq L$ and $\max(\mathbf{B}^{t})\geq U$, which is a contradiction. Therefore, $U=L$ and the model converges to consensus.